the question reads : evaluate expression for the given value of the variable

2x to the second power minus 5x+3, if x=1

Answer by stanbon(75887)

evaluate expression for the given value of the variable
2x to the second power minus 5x+3, if x=1
------
f(x) = 2x^2-5x+3
----
f(1) = 2*1-5*1+3 = 0
-----------------
Cheers,
Stan H
-------------